OpenStudy (anonymous):

Find the integral of (tan(theta)-1)^2 d*theta.

OpenStudy (anonymous):

first expand the bracket: \[\int\limits_{}^{}(\tan(\theta) - 1)^2 d\theta = \int\limits_{}^{}[ \tan^2(\theta) - 2\tan(\theta) + 1] d\theta\]use the trig id: sec^2(x) = tan^2(x) + 1 \[\int\limits_{}^{}[\sec^2(\theta) + 2\tan(\theta)]d\theta\] lmk if i need to keep going ^_^